calories to melt 17g of ice



1
Expert's answer
2021-03-04T06:32:37-0500

Heat of fusion (Lf) is the amount of energy in the form of heat needed to change the state of matter from a solid to a liquid (melting.)

Lf(H2O)=80cal/g;

Q-?


Q=m*Lf=17*80=1360 cal


Answer: 1360 calories are needed to melt 17 g of ice